# Prophet: Prompting Large Language Models with Complementary Answer Heuristics for Knowledge-based Visual Question Answering

### v3

@@ 4.1 Datasets
OK-VQA is a commonly used knowledge-based VQA dataset [7]. The dataset contains 9K and 5K image-question pairs for training and testing, respectively. All questions are manually filtered to ensure that outside knowledge is required to answer the questions. Each data sample is annotated with ten open-ended answers. The accuracy computed by the soft scores is used as the evaluation metric [47]. We use the 1.1 version of OK-VQA in the experiments.
A-OKVQA is currently the largest knowledge-based VQA dataset [8]. The dataset is split into three subsets: 17K training, 1K validation, and 7K testing. Each question is annotated with ten open-ended answers for direct answer (DA) evaluation. Besides, it provides a multiple choice (MC) evaluation to choose the correct answer from four choices.
ScienceQA is a dataset that consists of about 21K questions over a diverse set of science topics [23]. Out of the 21K questions, only the ‘IMG’ subset of 10.3K (48.7%) samples have image content, which is used in our experiments. Consequently, the retained dataset consists of 6.2K training, 2.1K validation, and 2.0K testing samples. The questions require high school-level science knowledge to arrive at the correct answer chosen from multiple choices.
TextVQA contains 28K images and 45K questions, where each question requires models to read and reason about the text in the image to give a correct answer [24]. The dataset is split into three subsets of 34.6K training, 5K validation, and 5.7K testing questions. Similar to OK-VQA, each question is annotated with ten open-ended answers by humans and soft-voting accuracy is used as the evaluation metric. Following the strategy in [48, 49], we supplement the training set with the augmented VQA samples from ST-VQA [50].
4.2 Implementation Details
@@
ViLBERT, ${}^{\dagger}$ 35.28 44.97 (+9.69)
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Settings on other datasets. The settings and strategies for OK-VQA can be directly transferred to A-OKVQA to address its DA task. For the MC task, we follow the strategy in [8] to project the predicted answer to the nearest answer choice. Moreover, we design a Prophet variant for the MC task. It uses a slightly different prompt by adding the multiple choices to in-context examples and testing input, and instructs the LLM to choose the correct one from four choices.
@@
Settings of other VQA models. In addition to MCAN, we also experiment with one generative VQA model mPLUG [19], which is first pretrained on task-agnostic image-text corpus and then finetuned on specific VQA dataset. Following the aforementioned two-stage transfer learning paradigm for MCAN, the pretrained mPLUG model is first finetuned on the VQAv2 dataset and then further finetuned on specific knowledge-based VQA dataset.
@@ 4.3 Ablation Studies
We conduct ablation experiments for Prophet on OK-VQA using the default settings above. Results shown in Table I and Fig. 4 are discussed in detail below.
Prompting vs. retrieval. Prophet uses a prompting-based paradigm to predict the answer based on a set of promising answer candidates. In contrast, a previous work MAVEx [12] exploits answer candidates but adopts a retrieval-based paradigm to search knowledge from external KBs to determine the answer. As both Prophet and MAVEx train a VQA model to generate answer candidates (stage-1), we can compare the superiority of the two paradigms (stage-2). In Table I, we show the performance of the two paradigms in terms of stage-1 accuracy and final accuracy, respectively.
@@
Capability of VQA models. In Table I, we study how the VQA models of different capabilities impact the performance of Prophet. To better control the model capability, we use the same MCAN model trained with four visual features: region-based Bottom-Up [25] and VinVL [26] features and grid-based CLIP features from two backbones (ViT-L/14 and RN50$\times$64) [51]. Results show that more powerful VQA models (reflected in the stage-1 accuracies) lead to better performance of Prophet, as they provide answer heuristics of higher quality. Combining the results in Table I, we also observe that more powerful VQA models achieve less relative improvements from GPT-3, which can be explained by the intrinsic diminishing return property. As a by-product, we verify that the visual features are important to the performance of knowledge-based VQA, which is consistent with the observations in [17]. The models with CLIP-based visual features significantly outperform those with region-based features, indicating that the CLIP’s visual features contain richer visual knowledge due to large-scale pretraining.
[…]
Answer candidates. Table I varies the number of answer candidates $K$ from 0 to 10 to explore its effect on Prophet. For each testing sample, if the ground-truth answer is hit by one of the $K$ answer candidates, we accumulate the soft score of that ground-truth answer [^4] In practice, multiple ground-truth answers are provided. If multiple answers are hit simultaneously, we choose the answer with the largest soft score for accumulation.. The hit rate is calculated over the testing set by dividing the accumulated score by the number of samples.
@@
Example selection strategy. To show the effectiveness of our answer-aware example selection strategy, we compare it to other example selection strategies in Table I. The compared strategies include: (a) rand: examples that are randomly selected; (b) ques + img: examples that are selected based on the joint similarity of question and image features, which is used in PICa; (c) fused: our default strategy that selects examples based on the similarity of fused features; (d) fused + ques + img: a combination of our default strategy and PICa’s strategy; and (e) answer logits: examples that are selected based on the similarity of answer logits obtained in Eq.(3). Besides the final accuracy, we also report the hit rate of answers within the selected examples for each strategy.
@@
Numbers of examples and queries. Table I contains the ablation studies for the numbers of examples and queries. We choose different numbers of examples $N\in\{0,1,8,16,20\}$ for each query and different numbers of queries $T\in\{1,5\}$, respectively. The results show that the performance of Prophet improves with the increase of $N$ and $T$, which is consistent with the results in PICa. By increasing $T$ from 1 to 5, the entries with larger $N$ enjoy greater performance improvements at the expense of linearly increasing overheads.
@@
Prompt contents. In Table I, we ablate the prompt contents in the default settings by: (b) removing the prompt head; (c) removing the confidence scores for answer candidates; (d) removing image captions; and (e) adding predicted tags from external models [15].
@@
Figure 4: Prophet’s prediction behaviors in terms of (a) distribution and (b) per-type accuracy. As Prophet takes $K$ answer candidates as inputs, we define three prediction behaviors for Prophet as follows: “keep top-1”, “in top 2-$K$”, and “beyond top $K$”. All the testing samples can be categorized into one of the three classes.
Stage 1 pred. Stage 2 pred.     correct     wrong
correct 54.4% 4.2%
wrong 12.0% 29.4%
TABLE II: Prophet’s combinatorial prediction behaviors in two stages. Prophet maintains the majority of correct predictions at stage-1, and the accuracy improvement by stage-2 is mainly because the number of wrong-to-correct samples is larger than that of the correct-to-wrong samples.
Prediction behaviors in different stages. In Table I, we can observe a significant performance improvement of Prophet (stage-2) over its corresponding MCAN model (stage-1). To better understand this improvement, we conduct a statistical analysis of Prophet’s prediction behaviors. As Prophet takes $K$ answer candidates from MCAN as inputs, we define three prediction behaviors for Prophet: “keep top-1”, “in top 2-$K$”, and “beyond top-$K$”. All the testing samples can be categorized into one of the three classes. The statistical results in Figure 4 show that: (i) for 68.1% of the testing samples (green slice), Prophet keeps the top-1 predictions of MCAN. These samples achieve a 69% accuracy and are mostly easy samples; (ii) for 21.8% of the testing samples (blue slice), Prophet selects answers from the top 2-$K$ answer candidates. These samples are relatively hard, so that MCAN delivers a 24% accuracy while Prophet has a much higher 40% accuracy; (iii) for the remaining 10.1% of the testing samples (yellow slice), Prophet predicts answers beyond the answer candidates [^5] The probability that Prophet’s prediction is constituted of the combination of candidates is rare that can be neglected.. For these most difficult samples, MCAN only delivers a 12% accuracy while Prophet magnificently achieves a 42% accuracy.
As a supplement to the above results, we calculate the distribution of four situations of the predictions from stage-1 and stage-2 in Table II. From the results, we can see that: (i) Prophet maintains the majority of correct predictions by MCAN and only 4.2% samples are overturned; (ii) the improvement of Prophet is mainly due to the fact that the proportion of wrong-to-correct samples (12.4%) is larger than that of the correct-to-wrong samples (4.2%); (iii) there are still a considerable amount of samples (29.4%) that both MCAN and Prophet fail to give the correct answer, which leaves sufficient room for future improvement.
@@
${}^{{\dagger}}$aMA-1 (7B) [20] 2.6s 51.8
${}^{{\dagger}}$aMA-1 (13B) [20] 4.6s 56.1
${}^{{\dagger}}$aMA-1 (30B) 8.7s 57.1
${}^{{\dagger}}$aMA-1 (65B) [20] 16.6s 58.8
${}^{{\dagger}}$alcon (7B) [21] 2.7s 50.5
${}^{{\dagger}}$alcon (40B) [21] 14.1s 57.1
@@
TABLE III: Ablation study of different LLMs. All variants use the default settings and are evaluated on the testing set of OK-VQA. The per-sample average costs of the open-source models are measured by the GPU running time on a server with A100 GPUs while the costs of the commercial models are measured by money. ${}^{{\dagger}}$ indicates the LLM’s max token length is insufficient for $N$=16 examples. For these LLMs, we reduce $N$ to fit their maximum capacity.
Different LLMs. In Table III, we investigate the effects of different LLMs by replacing the default GPT-3 (text-davinci-002) with the latest commercial and open-source models. From the results, we have the following observations: (i) the capability of the default GPT-3 model significantly outperforms all the compared LLMs, including its accelerated variant (3.5-turbo-instruct) with 0.075$\times$ running cost; (ii) for the LLMs of the same class but different sizes (e.g., 7B and 13B LLaMA-1 models [20]), the large-size ones show better performance than the small-size ones at the expense of near-linearly increasing running time; (iii) the chat-oriented variants like LLaMA-2-Chat [53], which are additionally trained by instruction tuning and human feedback [55], deliver inferior performance to their non-chatty counterparts. This can be explained by the introduced alignment tax when aligning the model with human behaviors; (iv) with only 7B model parameters, the latest LLM Mistral [54] reports near GPT-3 level performance, revealing the potential of open-source LLMs in the near future.
@@ 4.4 Main Results
For the comparisons below, we use all the default settings except the number of examples $N$. We set $N$=20 for OK-VQA and A-OKVQA and respectively set $N$=7 and $N$=16 for ScienceQA and TextVQA as they need extra hint and OCR tokens. By instantiating Prophet with two VQA models, we obtain Prophet (MCAN) and Prophet (mPLUG).
Comparative results on OK-VQA. Table IV contains the comparisons of our Prophet and existing state-of-the-art methods on OK-VQA. The table is split into three sections. The first section lists the retrieval-based methods leveraging external KBs [10, 56, 11, 42, 12, 13]. The second section contains the methods that are directly pretrained on a large-scale multimodal corpus [57, 2, 58]. The last section shows the methods that incorporate the large language model GPT-3, which is publicly available via an online API [15, 16, 17, 43].
@@
${}^{\dagger}$ [16] 53.1
${}^{\dagger}$ [17] 56.6
PromptCap (${}^{\dagger}$ [43] 60.4
@@
TABLE IV: Comparisons to the state-of-the-art methods on OK-VQA testing set. The compared methods are split into three groups based on their knowledge resources and usages. $*{}^{*}$: accuracy is evaluated on OK-VQA v1.0. ${}^{\dagger}$: method needs to query GPT-3 during training.
Our Prophet belongs to the last section. It outperforms all the compared methods by a distinct margin. Prophet is 13.1 points higher than PICa [15] when both methods use GPT-3 as the only knowledge resource. This confirms our hypothesis that the capacity of GPT-3 has not been fully activated in previous studies. Compared to KAT [16] and REVIVE [17], which utilize GPT-3 and other external KBs together in sophisticated systems, our Prophet is much simpler and more effective. Moreover, KAT, REVIVE, and PromptCap need to use GPT-3 to process all the training samples for their model training, which significantly increases the costs. In contrast, our Prophet only uses GPT-3 at inference time, which is more economical. Compared to the Flamingo-80B equipped with 32 in-context examples [2], Prophet (MCAN) delivers a significant performance improvement. Despite the fact that Prophet (MCAN) has a clear performance gap compared to PALI-17B [58], Prophet is more resource-efficient from the perspective of reproducibility [^6] Flamingo-80B is trained on 1,536 TPUv4 for 15 days and PALI is trained on 1,024 TPUv4 for 7 days, which are unaffordable for most researchers. In contrast, Prophet (MCAN) uses one RTX-3090 to train a VQA model for 4 days and a certain number of GPT-3 invocations.. Finally, by replacing MCAN with the pretrained generative model mPLUG, our method exhibits a 1.4-point further improvement, showing the substantial contribution of a powerful VQA model for Prophet.
@@
Comparative results on A-OKVQA. Table V contains the comparative results on the challenging A-OKVQA dataset. The results on the DA task show that Prophet (MCAN) model significantly outperform most existing approaches, reflecting the effectiveness and generalization of our method. Compared to the current state-of-the-art method PromptCap [43] which also involves a pretrained VQA model OFA [37] and GPT-3, Prophet (MCAN) exhibits similar performance when using a weaker VQA model.
For the MC task, we introduce a Prophet variant with slightly modifying the prompt used in the original Prophet. In particular, we add the multiple-choice information into both the in-context examples and testing input to instruct GPT-3 to choose the correct one from four choices. Compared with all the methods, Prophet (MCAN) surpasses all the counterparts on the MC task, showing the flexibility and scalability of Prophet. Moreover, Prophet (mPLUG) steadily outperforms Prophet (MCAN), emphasizing the significance of a powerful VQA model to Prophet.
Results on ScienceQA and TextVQA To verify the generalization ability of Prophet, we conduct experiments on two additional knowledge-based VQA datasets ScienceQA (IMG) and TextVQA, which require different types of knowledge (i.e., scientific knowledge and OCR knowledge) than that for OK-VQA and A-OKVQA. A Table VI shows that comparative results of Prophet and existing state-of-the-art methods on respective datasets. As we have witnessed the steady improvements of mPLUG over MCAN, we only report the results for Prophet (mPLUG) on these two datasets. Specifically, Prophet surpasses all the counterparts on ScienceQA (IMG), including the average human performance [23] and the latest LLaVa model trained with visual instruction tuning [60]. On TextVQA, Prophet outperforms the published state-of-the-art methods, including those methods with text-aware or layout-aware pretraining on large-scale scene-text image datasets [49, 48].
method accuracy
@@
[…]
method accuracy
@@
[…]
TABLE VI: Comparisons to the state-of-the-art methods on the testing set of ScienceQA and TextVQA, respectively.
4.5 Qualitative Analysis
5 Broader Impact
From a multimodal LLM (MLLM) point of view, Prophet is a loosely-coupled MLLM consisting of a vision-language (VL) model and a frozen LLM, aiming to endow the VL model with knowledge reasoning ability. Compared with the tightly-coupled MLLMs (e.g., Flamingo [2] and LLaVa [60]) which jointly optimize the VL model and LLM in an end-to-end manner, Prophet is more flexible that can support any open-source or commercial LLM.
Moreover, Prophet can also be regarded as a learning-to-prompt paradigm that learns an external model to generate prompts to better comprehend the target task, thus facilitating the capability of the pretrained LLM (or MLLM). From this point of view, recent studies like VoxPoser [68] and SoM-Prompting [69] share a similar idea with our work. We believe this paradigm can be widely used in a variety of LLM-related tasks.

### v4

@@ 4.1 Datasets
OK-VQA is a commonly used knowledge-based VQA dataset [15]. The dataset contains 9K and 5K image-question pairs for training and testing, respectively. All questions are manually filtered to ensure that outside knowledge is required to answer the questions.
A-OKVQA is currently the largest knowledge-based VQA dataset [16]. The dataset is split into three subsets: 17K training, 1K validation, and 7K testing. Both direct answering (DA) and multiple choice (MC) evaluations are provided.
ScienceQA is a dataset that consists of about 21K questions over a diverse set of high school-level science topics [25]. Out of the 21K questions, only the ‘IMG’ subset of 10.3K (48.7%) samples with images are used in our experiments.
TextVQA contains 28K images and 45K questions, where each question requires models to read and reason about the text in the image [26]. The dataset is split into three subsets of 34.6K training, 5K validation, and 5.7K testing questions. Following [43, 44], we supplement the training set with the augmented VQA samples from ST-VQA [45].
@@
ViLBERT, prompt 35.28 44.97 (+9.69)
(a) Prompting vs. retrieval. Our prompting-based paradigm is more effective than the retrieval-based one in MAVEx [19]. : our re-implementation.
@@
(b) Capability of VQA models. More powerful VQA models lead to higher accuracies, but obtain slightly less relative improvements from stage-2.
@@
(c) Answer candidates. They are crucial to Prophet and increasing $K$ leads to better performance.
@@
(d) Example selection strategy. Our answer-aware example selection based on fused features is more effective than the others.
@@
(e) Numbers of examples and queries. Increasing $N$ and $T$ improves model performance at the expense of linearly increasing overheads.
@@
(f) Prompt contents. The default settings contain the exact necessary information for prompting.
@@
4.2 Implementation Details
Settings on other datasets. The settings and strategies for OK-VQA can be directly transferred to A-OKVQA to address its DA task. For the MC task, we follow the strategy in [16] to project the predicted answer to the nearest answer choice. Moreover, we design a Prophet variant for the MC task. It uses a slightly different prompt by injecting the choices to in-context examples and testing input, and instructs the LLM to choose the correct one from four choices.
@@ 4.2 Implementation Details
Settings for other Prophet and Prophet++ variants. In addition to MCAN, we also experiment on Prophet with a generative VQA model mPLUG [40], which is first pretrained on massive image-text pairs and then finetuned on specific VQA dataset. Following the aforementioned two-stage transfer learning paradigm for MCAN, the pretrained mPLUG model is first finetuned on the VQAv2 dataset and then further finetuned on specific knowledge-based VQA dataset. For Prophet++, we use the state-of-the-art LMM GPT-4o in both stages.
@@ 4.3 Ablation Studies
We conduct ablation experiments for Prophet and Prophet++ on OK-VQA using the default settings above if not mentioned otherwise. Results shown in Table I and Fig. 5 are discussed in detail below.
Prompting vs. retrieval. Prophet uses a prompting-based paradigm to predict the answer based on a set of promising answer candidates. In contrast, a previous work MAVEx [19] exploits answer candidates but adopts a retrieval-based paradigm to search knowledge from external KBs to determine the answer. As both Prophet and MAVEx train a VQA model to generate answer candidates (stage-1), we can compare the superiority of the two paradigms (stage-2). In Table I(a), we show the performance of the two paradigms in terms of stage-1 accuracy and final accuracy, respectively.
@@
Capability of VQA models. In Table I(b), we study how the VQA models of different capabilities impact the performance of Prophet. To better control the model capability, we use the same MCAN model trained with four visual features: region-based Bottom-Up [2] and VinVL [27] features and grid-based CLIP features from two backbones (ViT-L/14 and RN50$\times$64) [46]. Results show that more powerful VQA models (reflected in the stage-1 accuracies) lead to better performance of Prophet, as they provide answer heuristics of higher quality. Combining the results in Table I(a), we also observe that more powerful VQA models achieve less relative improvements from GPT-3, which can be explained by the intrinsic diminishing return property. As a by-product, we verify that the visual features are important to the performance of knowledge-based VQA, which is consistent with the observations in [23]. The models with CLIP-based visual features significantly outperform those with region-based features, indicating that the CLIP’s visual features contain richer visual knowledge due to large-scale pretraining.
In addition to using different visual features for MCAN, we can also replace the whole MCAN model with any generative models pretrained on large-scale multimodal datasets as mentioned in §$\S$3.1. These results will be reported in the main results.
Answer candidates. Table I(c) varies the number of answer candidates $K$ from 0 to 10 to explore its effect on Prophet. For each testing sample, if the ground-truth answer is hit by one of the $K$ answer candidates, we accumulate the soft score of that ground-truth answer [^4] In practice, multiple ground-truth answers are provided. If multiple answers are hit simultaneously, we choose the answer with the largest soft score for accumulation.. The hit rate is calculated over the testing set by dividing the accumulated score by the number of samples.
@@
Example selection strategy. To show the effectiveness of our answer-aware example selection strategy, we compare it to other example selection strategies in Table I(d). The compared strategies include: (a) rand: examples that are randomly selected; (b) ques + img: examples that are selected based on the joint similarity of question and image features, which is used in PICa; (c) fused: our default strategy that selects examples based on the similarity of fused features; (d) fused + ques + img: a combination of our default strategy and PICa’s strategy; and (e) answer logits: examples that are selected based on the similarity of answer logits obtained in Eq.(3). Besides the final accuracy, we also report the hit rate of answers within the selected examples for each strategy.
@@
Numbers of examples and queries. Table I(d) contains the ablation studies for the numbers of examples and queries. We choose different numbers of examples $N\in\{0,1,8,16,20\}$ for each query and different numbers of queries $T\in\{1,5\}$, respectively. The results show that the performance of Prophet improves with the increase of $N$ and $T$, which is consistent with the results in PICa. By increasing $T$ from 1 to 5, the entries with larger $N$ enjoy greater performance improvements at the expense of linearly increasing overheads.
@@
Prompt contents. In Table I(f), we ablate the prompt contents in the default settings by: (b) removing the prompt head; (c) removing the confidence scores for answer candidates; (d) removing image captions; and (e) adding predicted tags from external models [21].
@@
Figure 5: Prophet’s prediction behaviors in terms of (a) distribution and (b) per-type accuracy. As Prophet takes $K$ answer candidates from MCAN as inputs, we define three prediction behaviors of Prophet, namely “keep top 1”, “in top 2-$K$”, and “beyond top $K$” predictions of MCAN, respectively. Note that all the testing samples can be categorized into one of the three classes above.
Prediction behaviors in different stages. In Table I(b), we can observe a significant performance improvement of Prophet (stage-2) over its corresponding MCAN model (stage-1). To better understand this improvement, we conduct a statistical analysis of Prophet’s prediction behaviors. As Prophet takes $K$ answer candidates from MCAN as inputs, we define three prediction behaviors of Prophet, namely “keep top 1”, “in top 2-$K$”, and “beyond top $K$”, which means LLM’s prediction hits the top-1 candidate, one of the top 2-$K$ candidates, none of the $K$ candidates, respectively. Note that all the testing samples can be categorized into one of the three classes above. The statistical results in Fig. 5 show that: (i) for 68.1% of the testing samples (green slice), Prophet keeps the top-1 predictions of MCAN. These samples achieve a 69% accuracy and are mostly easy samples; (ii) for 21.8% of the testing samples (blue slice), Prophet selects answers from the top 2-$K$ answer candidates. These samples are relatively hard, so that MCAN delivers a 24% accuracy while Prophet has a much higher 40% accuracy; (iii) for the remaining 10.1% of the testing samples (yellow slice), Prophet resorts to the LLM to predict new answers beyond the answer candidates [^5] The probability that Prophet’s prediction is constituted of the combination of candidates is rare that can be neglected.. For these challenging samples, MCAN only delivers a 12% accuracy while Prophet magnificently achieves a 42% accuracy. More analyses are provided in the appendix.
@@
TABLE II: Ablation study of different LLMs. All variants use the default settings and are evaluated on the testing set of OK-VQA. The per-sample average costs of the open-source models are measured by the GPU running time on a server with A100 GPUs while the costs of the commercial models are measured by money.
method stage-1 stage-2 accu.
VQA model LMM LLM/ LMM pure text/ img+text
GPT-4o - - - - 46.0
Prophet MCAN - GPT-4o pure-text 62.9
Prophet++ MCAN GPT-4o GPT-4o pure-text 63.8
MCAN GPT-4o GPT-4o img+text 64.2
TABLE III: Ablation study for Prophet++ on the testing set of OK-VQA. The first split contains the results of direct-prediction and Prophet-based prediction using the same LMM (GPT-4o). The second split contains two Prophet++ variants with different settings. $N$=16 is used for Prophet and Prophet++.
Different LLMs. In Table II, we investigate the effects of different LLMs by replacing the default GPT-3 (text-davinci-002) with the latest commercial and open-source models. From the results, we have the following observations: (i) the capability of the default GPT-3 model consistently outperforms all the compared LLMs, including its accelerated variant (3.5-turbo-instruct); (ii) for the LLMs of the same version but different sizes (e.g., 7B and 13B LLaMA-2 models [49]), the large-size ones show better performance than the small-size ones at the expense of near-linearly increasing running time; (iii) the chat-oriented variants like LLaMA-2-Chat [49], which are additionally trained by instruction tuning and human feedback [51], deliver inferior performance to their non-chatty counterparts. This can be explained by the introduced alignment tax when aligning the model with human behaviors by RLHF; (iv) with only 7B model size, the latest LLM Mistral [50] reports near GPT-3 level performance, revealing the potential of open-source LLMs.
Effect of LMMs in Prophet++. In Table III, we investigate the effects of different strategies. The results suggest that: (i) although GPT-4o is a highly-capable LMM, directly using it in a direct-prediction manner achieves much worse performance than the Prophet-based counterpart (46.0 vs. 62.9), showing the effectiveness of Prophet’s two-stage prompting paradigm; (ii) by introducing the third type of answer heuristic in the first stage, Prophet++ delivers a 0.9-point improvement over Prophet, showing the effectiveness and complementarity of the answer-aware rationales; (iii) when visual signals are additionally utilized in the second stage, we obtain further a 0.4-point improvement, validating the effectiveness of multimodal prompting in Prophet++. More results with the latest LMMs are provided in the appendix.
@@ 4.4 Main Results
For the comparisons below, we use all the default settings except the number of examples $N$. We set $N$=20 for OK-VQA and A-OKVQA and respectively set $N$=7 and $N$=16 for ScienceQA and TextVQA as they need extra hint and OCR tokens. By instantiating Prophet with two VQA models, we obtain Prophet (MCAN) and Prophet (mPLUG). Finally, we report the result of Prophet++ (mPLUG) on all the datasets to compare with Prophet (mPLUG).
Comparative results on OK-VQA. Table IV contains the comparisons of our Prophet and existing state-of-the-art methods on OK-VQA. The table is split into three sections. The first section lists the retrieval-based methods leveraging external KBs [18, 19, 20]. The second section contains the methods that are directly pretrained on a large-scale multimodal corpus [52, 11, 53]. The last two sections show the methods that incorporate the GPT-3 and GPT-4o [21, 22, 23, 38], which are accessible via online APIs.
@@
KAT [22] 53.1
REVIVE [23] 56.6
PromptCap (OFA) [38] 60.4
@@
methods with GPT-4o API
Prophet++ (mPLUG) 65.7
TABLE IV: Comparisons to the state-of-the-art methods on OK-VQA testing set. The compared methods are split into three groups based on their knowledge resources and usages. : method needs to query GPT-3 during training.
Our Prophet belongs to the last section. It outperforms all the compared methods by a distinct margin. Prophet is 13.1 points higher than PICa [21] when both methods use GPT-3 as the only knowledge resource. This confirms our hypothesis that the capacity of GPT-3 has not been fully activated in previous studies. Compared to KAT [22] and REVIVE [23], which utilize GPT-3 and other external KBs together in sophisticated systems, our Prophet is much simpler and more effective. Moreover, KAT, REVIVE, and PromptCap need to use GPT-3 to process all the training samples for their model training, which significantly increases the costs. In contrast, our Prophet only uses GPT-3 at inference time, which is more economical. Compared to the Flamingo-80B equipped with 32 in-context examples [11], Prophet (MCAN) delivers a significant performance improvement. Despite the fact that Prophet (MCAN) has a clear performance gap compared to PALI-17B [53], Prophet is more resource-efficient from the perspective of reproducibility [^6] Flamingo-80B is trained on 1,536 TPUv4 for 15 days and PALI is trained on 1,024 TPUv4 for 7 days, which are unaffordable for most researchers. In contrast, Prophet (MCAN) uses one RTX-3090 to train a VQA model for 4 days and a certain number of GPT-3 invocations.. Moreover, by replacing MCAN with the pretrained generative model mPLUG, our method exhibits a 1.4-point further improvement, showing the substantial contribution of a powerful VQA model for Prophet. Finally, Prophet++ (mPLUG) equipped with a modern LMM delivers a 3.2-point improvement over Prophet (mPLUG).
@@
method accu.
@@
Prophet++ (mPLUG) 90.5
(a) ScienceQA (IMG)
method accu.
@@
Prophet++ (mPLUG) 61.8
(b) TextVQA
TABLE VI: Comparisons to the state-of-the-art methods on ScienceQA and TextVQA.
Comparative results on A-OKVQA. Table V contains the comparative results on the challenging A-OKVQA dataset. The results on the DA task show that Prophet (MCAN) model significantly outperform the existing approaches including its base VQA model MCAN, showing its effectiveness and generalizability. Compared to the state-of-the-art method PromptCap [38] which also involves a pretrained VQA model OFA [12] and GPT-3, Prophet (MCAN) exhibits similar performance when using a weaker VQA model.
For the MC task, we introduce a Prophet variant with slightly modifying the prompt used in the original Prophet. In particular, we add the multiple-choice information into both the in-context examples and testing input to instruct GPT-3 to choose the correct one from four choices. Compared with all the methods, Prophet (MCAN) surpasses all the counterparts on the MC task, showing the flexibility and scalability of Prophet. Moreover, by introducing a more powerful VQA model mPLUG, Prophet (mPLUG) consistently outperforms Prophet (MCAN), while Prophet++ (mPLUG) further outperforms Prophet (mPLUG).
Results on ScienceQA and TextVQA To verify the generalization ability of Prophet, we conduct experiments on two additional knowledge-based VQA datasets ScienceQA (IMG) and TextVQA, which require different types of knowledge (i.e., scientific knowledge and OCR knowledge) than that for OK-VQA and A-OKVQA. A Table VI shows that comparative results of Prophet and existing state-of-the-art methods on respective datasets. As we have witnessed the steady improvements of mPLUG over MCAN, we only report the results for Prophet (mPLUG) on these two datasets. Specifically, Prophet surpasses all the counterparts on ScienceQA (IMG), including the base VQA model mPLUG, the average human performance [25], and the latest LMM LLaVa trained with visual instruction tuning [42]. On TextVQA, Prophet outperforms the published state-of-the-art methods, including the methods with text-aware or layout-aware pretraining on large-scale scene-text image datasets [44, 43]. Finally, the Prophet++ models prominently outperform their corresponding Prophet counterparts on respective datasets.